Again, we had good beam all night and the ABS was stable.  Tancredi 
modified the trigger file physics.settings and this was implemented at 
01:05 after run number 791.  Adam modified the trip reset button on the 
HV GUI at the same time.  Compton was showing low polarization (11-16%), 
and could be due to the change in the run plan.  We flipped the beam 
helicity once per run, rather than once per fill.
At 08:25 the BLAST power supply tripped at the end of run number 803.
We took the run numbers 792, 795, 797-798, 801-802 (250K events), updated 
the RUNLIST and crunched the data with lrn.  
Yuan and Ben.
This archive was generated by hypermail 2.1.2 : Mon Feb 24 2014 - 14:07:29 EST